UW CHEMISTRY NMR

Announcments & Equipment Logs

August, 2014

EARLY ALERT : POWER SHUTDOWN NEXT WEEK – Friday, Sep. 5th

Users might have noticed the power disruption announced for Bagley Hall on Sep. 6th, Saturday.  This planned event will affect the NMR lab Rm:37 from the early morning of Saturday 6th until about 3.00 pm on the same day.   The facility instruments, therefore will be shutdown by the afternoon of Friday the 5th Sep.  and […]

Posted in Hardware, Spectrometer | No Comments »

AV 700 UPDATE: Returned to use with TXI probe

The problems with the av700 pneumatics board have been resolved. The AV700 is available for use with the solution TXI probe. The av700 will be changing to solids in a couple weeks.

Posted in Spectrometer, Usage | No Comments »

DRX499 : UPDATE – 19F OBSERVE FULLY FUNCTIONAL

After creating a dataset of choice, users can simply enter the macro fluorine  to invoke the default 19F parameter set that includes  1H decoupling. Please remember to adjust O1P to your required offset to observe 19F signal.    Included below is a sample 19F spectrum.  Blue spectrum is with 1H decoupling and the Red spectrum is without 1H […]

Posted in Uncategorized | No Comments »

AV500 – UPDATE : EVERYTHING LOOKS GOOD

System is tested to work fine.  Here is a solvent suppression snapshot done on the dummy sample i.e. 1:1 H2O+D2O.

Posted in Uncategorized | No Comments »

AV700 – UPDATE : PNEUMATICS BOARD MALFUNCTION

Probably due to a voltage surge this weekend past,  the pneumatics control board on 700 appears compromised. No VT regulation or sample transport is possible due to this and we are energetically trying to resolve the issue.  Please look for further updates in this space. The machine is unusable until then.

Posted in Hardware, Spectrometer | No Comments »

DRX499 : UPDATE – OPERATIONAL WITH LIMITED FUNCTIONALITY

The spectrometer with the TXO probe is operational fully with 1H and 13C channels. The 19F is not yet ready for prime use as I write this and I will post an update once that issue is resolved as well. The instrument is open for regular use, with the above caveat.

Posted in Hardware, Spectrometer, Usage | No Comments »

UPDATE: AV300, AV301, AND AV500 AVAILABLE FOR USE

The errors being seen on the av300, av301, and av500 from the power fluctuations yesterday have been resolved. The spectrometers are now available for use. Please contact the facility managers and use the logbooks should you find further problems. Thanks.

Posted in Usage | No Comments »

Errors in ‘rga’ on av300, av300, and av500

It appears the storms this morning caused some fluctuations in the electrical and gas lines to the NMR spectometers leading to issues with the consoles, as is being reported to the facility by users using the error logbooks currently. For the time being, please consider the NMR spectrometers offline until an NMR Facility Manager can […]

Posted in Uncategorized | No Comments »

Item Categories